# Customizing Text-to-Image Diffusion with Camera Viewpoint Control

### v1

@@ Appendix A Results
Using predicted masks for background losses.   In our main paper, we calculated the background losses (Eqn. 11) using ground truth masks from the dataset during training. We show in Table 4 that replacing these with predicted masks results in similar performance. We use Detic [105] with Segment Anything [44] to predict the object mask given the object category, such as car and teddy bear.
@@
Comparison to LoRA.   We compare our method to the customization method LoRA [32, 76] on the text and image alignment metrics here. The CLIP and DINO scores for LoRA are $0.275$ and $0.468$, respectively. In comparison, our CLIP and DINO scores are $0.253$ and $0.481$, respectively. Though the CLIP score is marginally lower, our method performs better in preserving the target concept while allowing additional camera pose control for the custom object.
@@
Table 4: Results with predicted masks and variable views. Our method works similarly even when using predicted masks instead of ground truth masks for mask-based losses. When varying the number of views, the performance gradually drops w.r.t. camera pose accuracy and image alignment.
Appendix B Evaluation.
@@
Appendix C Implementation Details
C.1 Our Method
@@
During inference, we use an image guidance scale of $3.5$ and a text guidance scale of $7.5$ in Eqn. 13. All images for evaluation are generated with $50$ sampling steps using the default Euler scheduler [35]. With these settings, the wallclock runtime to generate one image given $8$ reference views is $10$ seconds. We cache the reference view features for each customized model. For comparison, it takes $6$ seconds to generate an image given no reference views.

### v2

@@ Appendix B More Results
Comparison with SDEdit+SDXL.   We add a comparison to SDEdit with the Stable Diffusion-XL model here, which we found to perform worse or on par with the Stable Diffusion-1.5 version of the model. Figure 12 shows the quantitative comparison.
Comparison to IP-Adapter + ControlNet.   Here, we compare against IP-Adapter [99] + depth ControlNet [107] for object viewpoint control. Given a target viewpoint, we render the image using a trained NeRF model of the object and use its predicted depth as a condition in depth ControlNet. We select the recommended hyperparameters that also worked best qualitatively, i.e., $0.6$ scale for the IP-Adapter and $0.5$ as ControlNet conditioning scale. The CLIP and DINO scores for the baseline are $0.251$ and $0.497$, respectively, similar to our method’s CLIP and DINO scores of $0.253$ and $0.481$. But in pairwise user-study comparison, our method is preferred by $60\pm 3\%$ and $75.67\pm 2.6\%$ for text-alignment and image-alignment. Qualitatively, we also observe that applying multiple conditionings, including depth and image, leads to lower text alignment as shown in Figure 17 ($5^{\text{th}}$ row).
Comparison to LoRA.   We compare our method to the customization method LoRA [34, 76] on the text and image alignment metrics here. The CLIP and DINO scores for LoRA are $0.275$ and $0.468$, respectively. In comparison, our CLIP and DINO scores are $0.253$ and $0.481$, respectively. Though the CLIP score is marginally lower, our method performs better in preserving the target concept while allowing additional object viewpoint control for the custom object.
@@
Table 4: Results with predicted masks and variable views. Our method works similarly even when using predicted masks instead of ground truth masks for mask-based losses. When varying the number of views, the performance gradually drops w.r.t. camera-pose accuracy and image alignment.
Appendix C Evaluation.
@@
Figure 11: Training with predicted camera viewpoints. We observe similar performance with predicted cameras using $50$ training images.
Appendix D Implementation Details
D.1 Our Method
@@
During inference, we use an image guidance scale of $3.5$ and a text guidance scale of $7.5$ in Eqn. 12. All images for evaluation are generated with $50$ sampling steps using the default Euler scheduler [37]. With these settings, the wallclock runtime to generate one image given $8$ reference views is $10$ seconds. We cache the reference view features for each customized model. For comparison, it takes $6$ seconds to generate an image given no reference views.
@@ D.2 Baselines
Appendix E Change log
v1:   Original draft.
v2:   Aditional comparison to 3D-editing based methods in Section 4.1 and more ablation experiments in Appendix A. Updated citations and writing.